Compute the probability, Mathematics

Assignment Help:

From past experience a machine is termed to be set up correctly on 90 percent of occasions.  If the machine is set up correctly then 95 percent of good parts are expected however if the machine is not set up correctly then the probability of a good part is only 30 percent.

On a specific day the machine is set up and the first component produced and found to be good.  What is the probability that the machine is set up appropriately.

Solution:

This is shown in the form of a probability tree or diagram as given below:

2000_Compute the Probability.png

 Whereas: CS - Correct Setting

IS - Incorrect Setting

GP - Good Product

BP - Bad Product

P(CSGP) = 0.9 × 0.95 = 0.855

P(CSBP) = 0.9 × 0.05 = 0.045

P(ISGP) = 0.1 × 0.3 =   0.03

P(ISBP) = 0.1 × 0.7 =    0.07

                                   1.00

- Probability of obtaining a good part (GP) = CSGP or ISGP

                                                            = CSGP + ISGP

                                                            = 0.855 + 0.03 = 0.885

Note: Good parts may be produced while the machine is correctly set up and when also it's incorrectly setup. In 1000 trials, 855 occasions while its correctly setup and good parts produced as CSGP and 30 occasions while its incorrectly setup and good parts produced as ISGP.

- Probability that the machine is correctly set up after getting a good part.

= number of favoural outcomes/total possible outcomes

=P (CSGP)/P(GP)

= 0.855/ 0.885

= 0.966

Or

= P (CS|GP)

= P (CSGP)/P(GP)

= 0.855/ 0.885

= 0.966


Related Discussions:- Compute the probability

How many baseball cards does peter now have, Peter purchased 14 latest base...

Peter purchased 14 latest baseball cards for his collection. This increased the size of his collection through 35%. How many baseball cards does Peter now have? First, you must

Promote products and services, please let us know above promote products an...

please let us know above promote products and services..i gave the assignment from my collage

Probability, There are 20 defective bulbs in a box of 100 bulbs.if 10bulbs ...

There are 20 defective bulbs in a box of 100 bulbs.if 10bulbs are choosen at random then what is the probability of there are just 3defective bulbs

Compound interest, you have RM5O,OOO to invest,and two fund that you''d li...

you have RM5O,OOO to invest,and two fund that you''d like to invest in.The You-Risk-It Fund yields 14% interest.The Extra-Dull Fund yields 6% interest.Besause of college financial-

Linear relations, determine the equation that represent the following lines...

determine the equation that represent the following lines be sure to define your variable and show all of your work

Organized list strategy, i can not figer out my homework it says "USE THE M...

i can not figer out my homework it says "USE THE MAKE AN ORGANIZED LIST STRATEGY,Medeline bikes 4 laps around her neighborhood 2 times a week.How many laps does she bike in 8 weeks

Implicit - explicit solution, It's easier to describe an explicit solution,...

It's easier to describe an explicit solution, in this case and then tell you what an implicit solution is not, and after that provide you an illustration to demonstrate you the dif

Write Your Message!

Captcha
Free Assignment Quote

Assured A++ Grade

Get guaranteed satisfaction & time on delivery in every assignment order you paid with us! We ensure premium quality solution document along with free turntin report!

All rights reserved! Copyrights ©2019-2020 ExpertsMind IT Educational Pvt Ltd